# Local Topological Constraints on Berry Curvature in Spin–Orbit Coupled BECs

### v4

@@ 4.2 Direct analysis for the deformed metrics $\varepsilon>0$
In an adapted gauge at $p$ we have $\Theta^{(\pm)}|_{p}=d\phi_{\pm}|_{p}$ and $\nabla^{LC(\varepsilon)}\operatorname{vol}_{\mathrm{BZ}}|_{p}=0$. Moreover,
$\nabla^{LC(\varepsilon)}\Theta^{(\pm)}|_{p}=d\Theta^{(\pm)}|_{p}=\varepsilon\,F^{(\pm)}|_{p}.$
Since $T_{\varepsilon}=F^{(+)}\wedge(d\phi_{+}+\varepsilon\pi^{*}A^{(+)})+F^{(-)}\wedge(d\phi_{-}+\varepsilon\pi^{*}A^{(-)})$, at the point $p$ we obtain
$\nabla^{LC(\varepsilon)}T_{\varepsilon}|_{p}=\operatorname{vol}_{\mathrm{BZ}}\wedge\bigl(c^{(+)}\nabla^{LC(\varepsilon)}(d\phi_{+})|_{p}+c^{(-)}\nabla^{LC(\varepsilon)}(d\phi_{-})|_{p}\bigr).$
A direct computation of the Levi‑Civita connection of $g_{\varepsilon}$ (see Appendix A) gives
$\nabla^{LC(\varepsilon)}d\phi_{\pm}|_{p}=\varepsilon\,c^{(\pm)}\,\omega,$
where $\omega$ is a fixed horizontal 2‑form with $\omega(X,Y)=1$. Consequently,
$\bigl(\nabla^{LC(\varepsilon)}T_{\varepsilon}\bigr)(X,Y,Z)=\bigl(c^{(+)}\varepsilon c^{(+)}+c^{(-)}\varepsilon c^{(-)}\bigr)=\varepsilon\,(c^{(+)2}+c^{(-)2})\neq 0.$
Raising indices yields an off‑diagonal operator $J^{(\varepsilon)}$. The factor $a_{+}c^{(+)}(1+\lambda^{2})$ appears when we express the result in terms of the cohomological normalization of $\xi_{0}$ (see Appendix B for details). Non‑vanishing of the constants guarantees injectivity. ∎
@@
For every $\varepsilon>0$,
@@
The operator $\Psi_{p}^{(\varepsilon)}(\xi_{0})$ belongs to the image of the curvature tensor $R^{C_{\varepsilon}}$ (it arises from the linear part $\nabla^{LC(\varepsilon)}T_{\varepsilon}$). By the Ambrose–Singer theorem, it lies in the holonomy algebra. Since $\Psi_{p}^{(\varepsilon)}$ is injective, it provides a non‑zero off‑diagonal element. Hence the off‑diagonal part of the holonomy algebra has dimension at least one. ∎
4.3 Semicontinuity and transfer to the physical metric
The lower bound for $\varepsilon>0$ together with the semicontinuity of the holonomy dimension allows us to obtain the same bound at the physical endpoint $\varepsilon=1$.
Lemma 4.6 (Lower semicontinuity of off‑diagonal holonomy dimension).
The function $\varepsilon\mapsto\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})$ is lower semicontinuous on $[0,1]$.
Proof.
The curvature tensor $R^{C_{\varepsilon}}$ depends smoothly on $\varepsilon$ and on the point. At a fixed $p\in M$, consider the linear map
$\Phi_{\varepsilon}:\bigwedge^{2}T_{p}M\longrightarrow\mathfrak{so}(T_{p}M),\qquad\Phi_{\varepsilon}(X\wedge Y)=R^{C_{\varepsilon}}(X,Y).$
The off‑diagonal holonomy algebra at $p$ is the Lie subalgebra generated by $\operatorname{im}\Phi_{\varepsilon}$. The dimension of a Lie algebra generated by a continuous family of linear maps is lower semicontinuous (the rank of a continuous matrix‑valued function is lower semicontinuous [15]). Taking the maximum over a finite set of points that generate the whole algebra preserves lower semicontinuity. ∎
@@ 4.3 Semicontinuity and transfer to the physical metric
By Corollary 4.5, for every $\varepsilon>0$ we have $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1$. Lower semicontinuity (Lemma 4.6) implies that the set
$S=\{\,\varepsilon\in[0,1]\mid\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1\,\}$
is closed. Since $S$ contains the interval $(0,1]$, it must also contain the endpoint $\varepsilon=1$. Hence $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{1}})\geq 1$. ∎
@@ 4.4 Geometric and physical interpretation
The inequality $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ has a direct geometric meaning: the curvature of the synthetic gauge connection cannot be made block‑diagonal with respect to the splitting $T_{p}M=\mathcal{H}_{p}\oplus\mathcal{V}_{p}$. At least one independent off‑diagonal curvature operator persists, mixing momentum directions with the phase directions $\phi_{\pm}$. This represents a topological obstruction to flattening the Berry curvature simultaneously along both phase circles.
@@
The bound remains non‑trivial even when the total Chern number $c_{1}^{(+)}+c_{1}^{(-)}=0$, as long as the individual curvatures are non‑zero. This demonstrates that the mixed‑tensor‑rank invariant captures a local topological constraint that persists when global topological charges cancel.
Remark 4.9 (Robustness).
The bound depends only on the cohomology class $\omega]$ (through the mixed rank) and on the generic condition $\nabla^{LC}T\neq 0$. Both are stable under small perturbations of the metric or the connections. Hence the obstruction is robust against experimental imperfections.
4.5 Summary of the argument
The logical structure of the proof is summarised as follows:
@@ 4.5 Summary of the argument
    Deformation family: Introduce $g_{\varepsilon}$ interpolating between the product metric $g_{0}$ and the physical KK metric $g_{M}$.
@@
    Product case ($\varepsilon=0$): Apply the original PT theorem to $g_{0}$; obtain the trivial bound $r^{\sharp}_{0}=0$.
@@
    Deformed cases ($\varepsilon>0$): Perform a direct pointwise computation in adapted gauge; show that the obstruction map $\Psi_{p}^{(\varepsilon)}$ is injective, yielding $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1$.
@@
    Semicontinuity: Use lower semicontinuity of the holonomy dimension to transfer the bound from $\varepsilon>0$ to the endpoint $\varepsilon=1$.
@@
    Physical conclusion: The physical connection $\nabla^{C}$ satisfies the same non‑trivial lower bound, certifying an irreducible topological coupling between momentum and phase degrees of freedom.
This strategy rigorously adapts the topological insight of the PT theorem to the Kaluza–Klein geometry of the SOC BEC, while avoiding the technical pitfalls of directly applying a theorem that requires a global product structure.
@@ 5 Topological Lower Bound and Its Physical Consequences
With the geometric construction of Section 2 and the cohomological data of Section 3 at hand, we can now synthesize these results to obtain the main physical consequence: a topological lower bound on the off-diagonal holonomy of the synthetic gauge connection in the SOC BEC model.
@@ 5.1 Statement of the bound
Our central result is expressed by the following inequality, which follows from the application of the PT theorem via the deformation-semicontinuity argument developed in Section 4:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}\;=\;r-\dim\mathcal{K},$ (5.1)
where $r=\operatorname{rank}_{\mathbb{R}}([T])$ is the mixed tensor rank of the torsion class, and $\mathcal{K}$ is the obstruction kernel associated with the metric. For the physical Kaluza–Klein metric $g_{M}$, we have established that $r^{\sharp}=1$, leading to a non-trivial lower bound.
@@ 5.2 Main theorem: topological lower bound for SOC BECs
Remark 5.1.
The PT lower bound is applied along the deformation family $(g_{\varepsilon},T_{\varepsilon})$ for every $\varepsilon>0$; no statement relies on the singular limit $\varepsilon=0$, and the resulting holonomy estimate is stable under deformation since it depends only on the mixed cohomology class.
@@
The proof follows directly from the deformation-semicontinuity argument carried out in Section 4. We recall the main steps.
1. Topological and geometric invariants. The mixed tensor rank of the torsion class is $r=1$ (Theorem 3.6), a purely topological invariant. For the product metric $g_{0}$, the obstruction kernel has dimension $\dim\mathcal{K}_{0}=1$, leading to a reduced rank $r^{\sharp}_{0}=0$ (Corollary 3.12). For the deformed metrics $g_{\varepsilon}$ with $\varepsilon>0$, the obstruction kernel vanishes, $\mathcal{K}_{\varepsilon}=0$, yielding a reduced rank $r^{\sharp}_{\varepsilon}=1$ (Corollary 3.15).
2. Applying the PT framework.
  • Product case ($\varepsilon=0$): The PT theorem applies directly to $(M,g_{0},\nabla^{C_{0}})$ and gives the (trivial) lower bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{0}})\geq r^{\sharp}_{0}=0$.
  • Deformed cases ($\varepsilon>0$): For each $\varepsilon>0$, the curvature analysis in adapted gauge (Theorem 4.4) shows that the obstruction map $\Psi_{p}^{(\varepsilon)}$ is injective, providing an explicit off-diagonal curvature operator. By the Ambrose–Singer theorem, this guarantees $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1$ (Corollary 4.5).
3. Semicontinuity and passage to the physical metric. The function $\varepsilon\mapsto\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})$ is lower semicontinuous (Lemma 4.6). Since the bound holds for all $\varepsilon>0$, lower semicontinuity implies it also holds at the endpoint $\varepsilon=1$, i.e., for the physical connection $\nabla^{C}=\nabla^{C_{1}}$. This establishes inequality (5.2). ∎

### v5

@@ 4.2 Direct curvature analysis for $\varepsilon>0$
We establish the bound by exhibiting a nonzero off-diagonal curvature operator at every point, using the curvature evaluated on mixed inputs (one horizontal, one vertical).
1. Setup in adapted gauge. Fix an arbitrary point $p\in M$ and choose adapted gauge where $A^{(\pm)}|_{p}=0$ (Lemma 2.19). At $p$, the orthonormal frame is
$e_{1}=\partial_{k_{x}},\quad e_{2}=\partial_{k_{y}}\quad(\text{horizontal, }\mathcal{H}_{p}),\qquad e_{3}=\partial_{\phi_{+}},\quad e_{4}=\partial_{\phi_{-}}\quad(\text{vertical, }\mathcal{V}_{p}),$
and the metric reduces to $g_{\varepsilon}|_{p}=dk_{x}^{2}+dk_{y}^{2}+d\phi_{+}^{2}+d\phi_{-}^{2}$.
2. Curvature decomposition for mixed inputs. The curvature of $\nabla^{C_{\varepsilon}}=\nabla^{\mathrm{LC}(\varepsilon)}+\tfrac{1}{2}T_{\varepsilon}$ decomposes as
$R^{C_{\varepsilon}}(X,Y)\;=\;R^{\mathrm{LC}(\varepsilon)}(X,Y)\;+\;\tfrac{1}{2}\bigl[(\nabla^{\mathrm{LC}(\varepsilon)}_{X}T_{\varepsilon})(Y,\cdot)-(\nabla^{\mathrm{LC}(\varepsilon)}_{Y}T_{\varepsilon})(X,\cdot)\bigr]^{\sharp}\;+\;\tfrac{1}{4}[T_{X},T_{Y}],$ (4.2)
where $T_{X}(Z):=T_{\varepsilon}(X,Z,\cdot)^{\sharp}$.
To detect off-diagonal holonomy, we evaluate (4.2) on the mixed pair $X=e_{1}\in\mathcal{H}_{p}$, $Y=e_{3}\in\mathcal{V}_{p}$ and extract the off-diagonal projection
$\pi_{\mathrm{off}}\!\bigl(R^{C_{\varepsilon}}(e_{1},e_{3})\bigr)\;\in\;\mathrm{Hom}(\mathcal{H}_{p},\mathcal{V}_{p})\oplus\mathrm{Hom}(\mathcal{V}_{p},\mathcal{H}_{p}),$
which maps horizontal directions to vertical ones and vice versa.
3. Torsion endomorphisms at $p$. At $p$ in adapted gauge, the torsion is $T_{\varepsilon}|_{p}=c^{(+)}e^{1}\!\wedge e^{2}\!\wedge e^{3}+c^{(-)}e^{1}\!\wedge e^{2}\!\wedge e^{4}$, which is independent of $\varepsilon$. The torsion endomorphism table (Lemma B.4 of Appendix B) gives, in particular:
$T_{e_{1}}(e_{2})$ $=c^{(+)}e_{3}+c^{(-)}e_{4}$ $(\mathcal{H}\to\mathcal{V}),$
$T_{e_{1}}(e_{3})$ $=-c^{(+)}e_{2}$ $(\mathcal{V}\to\mathcal{H}),$
$T_{e_{3}}(e_{1})$ $=c^{(+)}e_{2},\qquad T_{e_{3}}(e_{2})=-c^{(+)}e_{1}$ $(\mathcal{H}\to\mathcal{H}).$
4. Leading contribution: the quadratic torsion term. The commutator $T_{e_{1}},T_{e_{3}}]$ is computed entry-by-entry in Proposition B.6 of Appendix B:
$[T_{e_{1}},T_{e_{3}}](e_{1})$ $=(c^{(+)})^{2}\,e_{3}+c^{(+)}c^{(-)}\,e_{4}$ $(\mathcal{H}\to\mathcal{V}),$ (4.3)
$[T_{e_{1}},T_{e_{3}}](e_{3})$ $=-(c^{(+)})^{2}\,e_{1}$ $(\mathcal{V}\to\mathcal{H}),$ (4.4)
$[T_{e_{1}},T_{e_{3}}](e_{4})$ $=-c^{(+)}c^{(-)}\,e_{1}$ $(\mathcal{V}\to\mathcal{H}).$ (4.5)
Every nonzero entry is off-diagonal: it maps $\mathcal{H}\to\mathcal{V}$ or $\mathcal{V}\to\mathcal{H}$. In particular, (4.4) gives
$\pi_{\mathrm{off}}\!\bigl(\tfrac{1}{4}[T_{e_{1}},T_{e_{3}}]\bigr)(e_{3})=-\frac{(c^{(+)})^{2}}{4}\,e_{1}\;\neq\;0\quad\text{for all }c^{(+)}\neq 0.$ (4.6)
5. Subdominance of the remaining terms. The other two contributions to $R^{C_{\varepsilon}}(e_{1},e_{3})$ are:
  • Levi-Civita curvature: All curvature 2-forms $\Omega^{a}{}_{b}$ of $g_{\varepsilon}$ are proportional to $e^{1}\!\wedge e^{2}$ (Appendix A). Evaluating on the mixed pair $(e_{1},e_{3})$ gives zero: $R^{\mathrm{LC}(\varepsilon)}(e_{1},e_{3})=0$.
  • Torsion derivative: Since the torsion has constant coefficients in adapted gauge and the connection coefficients are $\mathcal{O}(\varepsilon)$, the covariant derivative $\nabla^{\mathrm{LC}(\varepsilon)}T_{\varepsilon}$ is $\mathcal{O}(\varepsilon)$.
Therefore:
$\pi_{\mathrm{off}}\!\bigl(R^{C_{\varepsilon}}(e_{1},e_{3})\bigr)\;=\;\pi_{\mathrm{off}}\!\bigl(\tfrac{1}{4}[T_{e_{1}},T_{e_{3}}]\bigr)\;+\;\mathcal{O}(\varepsilon).$ (4.7)
The leading term (4.6) is a nonzero constant independent of $\varepsilon$, while the correction is continuous and vanishes at $\varepsilon=0$. Hence the total is nonzero for all $\varepsilon\in[0,1]$.
6. Holonomy generation. By the Ambrose–Singer theorem, the holonomy algebra $\mathfrak{hol}_{p}(\nabla^{C_{\varepsilon}})$ contains all curvature endomorphisms $R^{C_{\varepsilon}}(X,Y)$ for $X,Y\in T_{p}M$. Since $\pi_{\mathrm{off}}(R^{C_{\varepsilon}}(e_{1},e_{3}))\neq 0$, there exists at least one nonzero element in the off-diagonal component
$\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C_{\varepsilon}}):=\mathfrak{hol}_{p}(\nabla^{C_{\varepsilon}})\cap\bigl(\mathrm{Hom}(\mathcal{H}_{p},\mathcal{V}_{p})\oplus\mathrm{Hom}(\mathcal{V}_{p},\mathcal{H}_{p})\bigr).$
By homogeneity of the model (Proposition B.2, Appendix B), the curvature components are independent of $p$, so the bound holds globally:
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C_{\varepsilon}})\;\geq\;1\qquad\text{for every }p\in M,\ \varepsilon>0,\ c^{(\pm)}\neq 0.$ (4.8)
@@
For every $\varepsilon>0$ and every point $p\in M$,
@@
This is the direct statement of Theorem 4.5. ∎
4.3 Application to the physical connection
@@
This follows immediately from Theorem 4.5 by setting $\varepsilon=1$. The physical connection $\nabla^{C}$ is precisely $\nabla^{C_{1}}$, which satisfies the inequality for every point in $M$. ∎
Remark 4.8 (Direct nature of the argument).
The proof of Theorem 4.7 establishes the bound through direct pointwise computation of the curvature as an explicit function of the SOC BEC model parameters $(c^{(+)},c^{(-)})$. The vanishing of the obstruction kernel for $\varepsilon>0$ (Lemma 2.16) provides the crucial geometric input that makes the bound non-trivial, contrasting sharply with the trivial bound at $\varepsilon=0$.
@@ 4.4 Geometric and physical interpretation
The inequality $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ has a direct geometric meaning: the curvature of the synthetic gauge connection cannot be made block-diagonal with respect to the splitting $T_{p}M=\mathcal{H}_{p}\oplus\mathcal{V}_{p}$. At least one independent off-diagonal curvature operator persists, mixing momentum directions (crystal momenta $k_{x},k_{y}$) with the phase directions ($\phi_{+},\phi_{-}$). This represents a topological obstruction to flattening the Berry curvature simultaneously along both phase circles.
Remark 4.9 (Topological locking of degrees of freedom).
The off-diagonal holonomy can be interpreted as a form of topological locking: the momentum and phase degrees of freedom are entangled by the mixed cohomology class $\omega]$ in a way that cannot be undone by smooth gauge transformations or metric deformations preserving the cohomology classes $F^{(\pm)}]$. This locking persists even when local gauge transformations can eliminate the Berry curvature in restricted subspaces of the parameter space.
@@
The bound remains non-trivial even when the total Chern number vanishes:
$c_{1}^{\mathrm{tot}}:=c_{1}^{(+)}+c_{1}^{(-)}=\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}\left(F^{(+)}+F^{(-)}\right)=0.$
As long as the individual curvatures are non-zero ($c^{(\pm)}\neq 0$), the mixed rank remains $r=1$ and the obstruction kernel vanishes for $\varepsilon>0$, yielding $r^{\sharp}=1$. This demonstrates that the mixed-tensor-rank invariant captures a local topological constraint that persists when global topological charges cancel. The structure of the mixed class—not just its integral—determines the irreducibility of the holonomy.
Remark 4.11 (Robustness under perturbations).
The bound depends only on:
@@
    The cohomology class $\omega]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]$, which is a topological invariant.
@@
    The condition that $c^{(\pm)}\neq 0$, which ensures non-vanishing curvature.
Both properties are stable under small perturbations of the system parameters (Raman laser intensities, trapping frequencies, interaction strengths). Hence, the topological obstruction certified by $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ is robust against experimental imperfections and represents a generic feature of the SOC BEC phase diagram.
Remark 4.12 (Relation to the cohomological lower bound of [12]).
The lower bound of [12] applies to cohomologically calibrated connections on Riemannian product manifolds and yields the invariant $r^{\sharp}=\min(r-\dim\mathcal{K},\,1)$, where $r$ is the mixed tensor rank and $\mathcal{K}$ is the obstruction kernel determined by parallel-form strata. In the product limit $\varepsilon=0$ of our model, the metric $g_{0}$ is a Riemannian product on $T^{2}_{\mathrm{BZ}}\times T^{2}_{\mathrm{fiber}}$ with both factors flat, so $\mathcal{P}_{1}(T^{2}_{\mathrm{fiber}})=\mathbb{R}^{2}$ and $r^{\sharp}=0$: the bound is trivial, consistently with the fact that the product connection has reducible holonomy.
For $\varepsilon>0$, the Kaluza–Klein metric $g_{\varepsilon}$ is no longer a Riemannian product, and the theorem of [12] does not apply directly. Nevertheless, the algebraic mechanism that drives the transition $r^{\sharp}=0\to r^{\sharp}=1$ in [12]—namely, the collapse of the parallel-form stratum $\mathcal{P}_{1}$ from $\mathbb{R}^{2}$ to a proper subspace—has an exact counterpart in our setting: Proposition LABEL:prop:parallel-drop shows that $\dim\mathcal{P}_{1}^{\varepsilon}$ drops from $2$ to $1$ as soon as $\varepsilon>0$. The direct curvature computation of Theorem 4.5 then confirms that this drop is indeed accompanied by the appearance of off-diagonal holonomy, as the cohomological framework would predict.
This suggests that the cohomological invariant $r^{\sharp}$ of [12] may capture the relevant obstruction beyond the Riemannian product setting, at least for Kaluza–Klein deformations of product metrics where the parallel-form strata remain computable. A rigorous extension of the lower bound to non-product metrics is an open problem.
4.5 Summary of the proof strategy
The logical structure of the proof can be summarized as follows:
  1. 1.
    Deformation family (Section 2): We construct a smooth one-parameter family of metrics $\{g_{\varepsilon}\}_{\varepsilon\in[0,1]}$ interpolating between the Riemannian product metric $g_{0}=g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$ and the physical Kaluza–Klein metric $g_{M}=g_{1}$.
  2. 2.
    Topological invariants (Section 3): We compute the mixed tensor rank $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})=1$ and the obstruction kernels:
    $\dim\mathcal{K}_{0}=1\quad(\varepsilon=0),\qquad\dim\mathcal{K}_{\varepsilon}=0\quad(\varepsilon>0).$
    This yields reduced ranks $r^{\sharp}_{0}=0$ (trivial) and $r^{\sharp}_{\varepsilon}=1$ (non-trivial).
@@ 4.5 Summary of the proof strategy
    Product case ($\varepsilon=0$): The PT theorem applies directly, giving the trivial bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{0}})\geq 0$.
@@
    Deformed cases ($\varepsilon>0$): Through explicit computation of the curvature $R^{C_{\varepsilon}}(X,Z)$ for mixed inputs $X\in\mathcal{H}_{p}$, $Z\in\mathcal{V}_{p}$ (Theorem 4.5, Appendix B), we show that the off-diagonal projection $\pi_{\mathrm{off}}(R^{C_{\varepsilon}}(X,Z))\neq 0$ at every point, establishing $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1$ for every $\varepsilon\in(0,1]$ and every $c^{(\pm)}\neq 0$.
@@
    Physical connection ($\varepsilon=1$): The bound holds for the physical connection $\nabla^{C}=\nabla^{C_{1}}$ as a direct consequence of the previous step.
This strategy rigorously applies the PT cohomological framework to the specific Kaluza–Klein geometry of the SOC BEC model by establishing that the essential topological constraint—encoded in $r^{\sharp}_{\varepsilon}=1$ for $\varepsilon>0$—translates into a geometric lower bound on off-diagonal holonomy through direct computation of the mixed-input curvature in the model’s parameter space.
@@ 5 Topological Lower Bound and Its Physical Consequences
With the geometric construction of Section 2 and the cohomological data of Section 3 at hand, we now synthesize these results to obtain the main physical consequence: a topological lower bound on the off-diagonal holonomy of the synthetic gauge connection in the SOC BEC model.
@@ 5.1 Statement of the bound
The central result is expressed by the following inequality, which follows from the uniform maximal rank of the obstruction map proved in Section 4:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}\;=\;\min(r-\dim\mathcal{K},1),$ (5.1)
where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ is the mixed tensor rank of the torsion class, and $\mathcal{K}$ is the obstruction kernel. For the physical Kaluza–Klein metric $g_{M}$, we have established that $r^{\sharp}=1$.
@@ 5.2 Main theorem: topological lower bound for SOC BECs
The proof follows from the uniform deformation analysis of Section 4:
  1. 1.
    Topological and geometric invariants: The mixed tensor rank is $r=1$ (Theorem 3.6). For $\varepsilon>0$ the obstruction kernel vanishes, $\mathcal{K}_{\varepsilon}=0$ (Lemma 2.16), giving reduced rank $r^{\sharp}_{\varepsilon}=1$.
  2. 2.
    Uniform bound for all $\varepsilon>0$: Theorem 4.5 proves that the obstruction map $\overline{\Psi}_{p}^{(\varepsilon)}$ has maximal rank $r^{\sharp}=1$ for every $\varepsilon\in(0,1]$ and every point $p$. By Ambrose–Singer, this yields $\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C_{\varepsilon}})\geq 1$ uniformly (Corollary 4.6).
  3. 3.
    Physical endpoint: Since $\varepsilon=1$ belongs to the interval $(0,1]$, the bound holds directly for the physical connection $\nabla^{C}$.
Remark 5.2 (Methodological scope).
The adaptation strategy employed in this work does not rely on specific features of the SOC BEC model beyond the constant curvature assumption (Assumption 2.3). The key ingredients are:
  1. 1.
    a smooth deformation family $\{g_{\varepsilon}\}_{\varepsilon\in[0,1]}$ interpolating between a Riemannian product metric ($\varepsilon=0$) and the physical metric ($\varepsilon=1$);
  2. 2.
    explicit computation of the parallel-form strata $P^{k}(M_{i},g_{\varepsilon})$ for the deformed metrics;
  3. 3.
    verification that the obstruction kernel $K_{\varepsilon}$ has reduced dimension for $\varepsilon>0$ compared to the product limit.
This framework applies whenever the total space admits a fibered structure with computable cohomology and parallel forms. Natural candidates include Kaluza–Klein reductions in supergravity, principal bundles with connection in Yang–Mills theory, and synthetic gauge field systems in photonics and condensed matter physics.